OpenStudy (anonymous):

I need help solving the rest of this problem. Suppose a parabola has vertex (-4, 7) and also passes through the point (-3, 8). Write the equation of the parabola in vertex form.

OpenStudy (anonymous):

This is what I have so far. I plugged in the vertex and the point it passes through. y = a(x - h)^2 + K 8 = a(-3 - 4)^2 + 7 8 = a(-7)^2 + 7 After this I am not sure what to do. I think you subtract 7 from both sides but I am not sure. Any help is appreciated.

OpenStudy (anonymous):

I have done that. Now I am trying to solve for a so I can then plug it back into the original equation.

OpenStudy (anonymous):

You should be able to plug in the x & y coordinates that the parabola passes through (not the vertex), then you'll be able to solve for a

OpenStudy (mertsj):

1. Replace h with -4, k with 7, x with -3, y with 8. 2. Solve that equation for a 3. Rewrite the equation replacing h with -4, k with 7, and a with the value from step 2 4. Walla!!! You are done.

OpenStudy (anonymous):

"walla" lol is it that you're confused how to solve for a once you've plugged in all your values?

OpenStudy (anonymous):

Ok, I plugged in all the values. 8 = a(-3 - 4)^2 + 7. 8 = a(-7)^2 + 7, what do I do after this? Do I subtract 7 from both sides?

OpenStudy (anonymous):

Yesss

OpenStudy (anonymous):

Okay, so after I subtract 7 from both sides I am left with 1 = a(-7)^2. What do I do after this. This is where I got stuck.

OpenStudy (anonymous):

What's 7^2? Divide both sides by that number :)

OpenStudy (anonymous):

sorry, 7^2 and (-7)^2 equal the same thing

OpenStudy (anonymous):

They were just lost on the order of operations, not how to do the problem in general :)

OpenStudy (anonymous):

I know how to write the equation out with the points I am given. I typed the equation out in my first post.

OpenStudy (anonymous):

Did you get the solution for a though?

OpenStudy (anonymous):

I am working on it, give me a sec.

OpenStudy (anonymous):

Oh alright lol

OpenStudy (anonymous):

The post directly under that showed me what I already know. A lot of people were posting and if I missed a post by you, that doesn't give you the right to be a douche. I am not good at math and insulting people on here is not how you go about things.

OpenStudy (anonymous):

^ lol you're doing well, don't worry. :)

OpenStudy (anonymous):

a = 1 right?

OpenStudy (anonymous):

Waittt, lemme show you So you had 1=a(-7)^2 So first (-7)^2 = 49 So, 1 = 49a Divide both sides by 49 and you've got your answer :)

OpenStudy (anonymous):

I had written the equation wrong. So the equation should've looked like 8 = a(-3 + 4)^2 +7. 8 = a(1)^2 + 7 Subtract 7 from each side and it becomes 1 = a(1)^2

OpenStudy (anonymous):

OH then yeah you're right! Sorry lol

OpenStudy (anonymous):

It's ok, thanks for both of your guys' help. I appreciate it.

OpenStudy (anonymous):

No prob, lemme know if you need more help :)

OpenStudy (anonymous):

Will do!